A car that travels at 1.3 km/hr for 2.0 hrs, at 6.5 for 1.0 hr and at 4.5 for half an hour. What is its average speed?


1
Expert's answer
2021-02-23T09:59:46-0500
"v_{av}=\frac{s}{t}\\v_{av}=\frac{2(1.3)+1(6.5)+0.5(4.5)}{2+1+0.5}=3.2\frac{km}{h}"